x-y+2z=-7
y+z=1
x=2y+3z
I did start this problem but got stuck fairly quickly. We're supposed to be solving using substitution or elimination. Please get back to me as soon as possible, thanks!

Answer by Alan3354(69443)

x-y+2z=-7
y+z=1
x=2y+3z
==============
Sub for x in the 1st eqn.
(2y + 3z) - y + 2z = 7
y + 4z = 7
y + z = 1
=================== Subtract
3z = 6
z = 2 *******************
------
y + z = 1
y = -1 **********************
------
Sub for y & z anywhere to find x.